\begin{thebibliography}{10}

\bibitem{AlHa2010}
A. Almeida, P. H\"{a}st\"{o}, Besov spaces with variable smoothness and integrability, J. Funct. Anal.,  258 (2010), 1628--1655.

\bibitem{GG}
A. Ghorbanalizadeh, P. Górka, Completeness and separability of the spaces of variable integrability and summability, Proc. Am. Math. Soc. 149 (09)(2021), 3873-3879

\bibitem{Cruz}
D. Cruz-Uribe, A. Fiorenza, Variable Lebesgue spaces: foundations and harmonic analysis, Springer Science  Business Media, 2013.

\bibitem{Nakai}
E. Nakai and Y. Sawano, Hardy spaces with variable exponents and generalized Campanato spaces, J. Funct. Anal.  262 (2012), no. 9, 3665--748.

\bibitem{LiLiPi}
F. Li, Z. Li and L. Pi, Ling, Variable exponent functionals in image restoration, Appl. Math. Comput. 216 (2010), no. 3, 870--882.

\bibitem{GRS}
Ghorbanalizadeh, A., Roohi Seraji, R. and Sawano, Y. On the reflexivity of the spaces of variable integrability and summability. Rev Mat Complut (2022). https://doi.org/10.1007/s13163-022-00447-w

\bibitem{GKP}
Gupta, M., Kamthan, P.K., Patterson, J.: Duals of generalized sequence spaces. J. Math. Anal. Appl. 82, 152–168 (1981)

\bibitem{KeVy}
H. Kempka, J. Vyb\'{i}ral,A note on the spaces of variable integrability and summability of Almeida and H\"{a}st\"{o}, Proc. Amer. Math. Soc.  141 (2013), 3207--3212.

\bibitem{Die}
L. Diening, P. Harjulehto, P. H\"{a}st\"{o}, M. R\u u\^zi\^cka, Lebesgue and Sobolev Spaces with Variable Exponents, Lecture Notes in Mathematics, Springer, Heidelberg, 2011.

\bibitem{di} 
L. Diening, P. H\"{a}st\"{o}, A. Nekvinda, Open problems in variable exponent Lebesgue and Sobolev spaces, in: Proc. Int. Conference Function Spaces, Differential Operators and Nonlinear Analysis, Milovy, Czech Rep., May 28 - June 4, 2004, Math. Inst. Acad. Sci. Czech Rep., Prague, 2005, 38-58.

\bibitem{LP}
Lin, Pei-Kee (1-MEMP) K\"{o}the-Bochner function spaces. Birkhäuser Boston, Inc., Boston, MA, 2004. xiv+370 pp. 

\bibitem{Day}
M. Day, Some more uniformly convex spaces, Bull. Amer. Math. Soc., {\bf 47} (1941), 504-507.

\bibitem{Ruzicka}
M. R\r{u}\^{z}i\^{c}ka, Electrorheological Fluids: Modeling and Mathematical Theory, Springer-Verlag, Berlin, 2000.

\bibitem{has}  
P. H\"{a}st\"{o}, \emph{Open problems}. presented at conference "Nonstandard growth phenomena", August 29-31, 2017, Turku, Finland.

\bibitem{Helly'sTheorem}
Robinson, J. (2020). An Introduction to Functional Analysis. Cambridge: Cambridge University Press. doi:10.1017/9781139030267

\bibitem{sam}
S. Samko, On a progress in the theory of Lebesgue spaces with variable exponent: maximal and singular operators, Integral Transforms Spec. Funct. 16 (2005), 461--482.

\bibitem{Zhikov2} 
V. V. Zhikov, Averaging of functionals in the calculus of variations and elasticity, Izv. Akad. Nauk SSSR Ser. Mat.  50 no. 4 (1986), 675--710.

\bibitem{WO}
W. Orlicz,   \"Uber konjugierte exponentenfolgen. Studia Mathematica 3, no. 1 (1931): 200--211.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{textcomp}
\usepackage{color}
\usepackage{graphicx}
\usepackage{pict2e}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{prop}[theorem]{Proposition}

\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}

\newcommand{\ci}{{\ell}^{q(\cdot)}(L^{p(\cdot)})}
\newcommand{\pe}{L^{p(\cdot)}(\Omega)}
\newcommand{\p}{{p( \cdot )}}
\newcommand{\q}{{q( \cdot )}}

\numberwithin{equation}{section}

\begin{document}

\title[Some Remarks on the Vector-Valued spaces $\ci$]{Some Remarks on the Vector-Valued Variable Exponent Lebesgue Spaces $\ci$}

\author{Arash Ghorbanalizadeh}
\address{Department of Mathematics, Institute for Advanced Studies in Basic Sciences (IASBS), Zanjan 45137-66731, Iran}
\curraddr{}
\email{ghorbanalizadeh@iasbs.ac.ir}

\author{Reza Roohi Seraji}
\address{
Department of Mathematics, Institute for Advanced Studies in Basic Sciences (IASBS), Zanjan 45137-66731, Iran}
\curraddr{}
\email{rroohi@iasbs.ac.ir}

\subjclass[2010]{Primary 46E30, Secondary 42B35}

\keywords{variable mixed Lebesgue-sequence space, strict convexity, uniform convexity, measure convergence, approximation identity}

\date{}

\dedicatory{}

\commby{}


\maketitle

\begin{abstract}
In this paper, we investigate the geometric properties of the variable mixed Lebesgue-sequence space  $\ell^{q(\cdot)} (L^{p(\cdot)})$  as a Banach space. We show that, if $ 1<q_-,p_-,q_+,p_+<\infty $, then $\ell^{q(\cdot)} (L^{p(\cdot)})$ is strictly and uniformly convex. We also prove that, when $ 1\le q_-,p_-,q_+,p_+<\infty, $  the convergence in norm implies the convergence in measure, and under some conditions on exponents, the approximation identity holds in the space $ \ell^1(L^{\p\slash\q}) $.
\end{abstract}

\section{introduction}
Variable Lebesgue spaces are a type of function spaces that generalize the classical Lebesgue spaces by allowing the exponent $p$ to vary with the point $x$. These spaces were originally introduced by Orlicz \cite{WO} and are denoted by $L^{p(\cdot)}$. They have many properties in common with Lebesgue spaces, but they also exhibit some distinctive and subtle features. Variable Lebesgue spaces are not only interesting from a mathematical point of view, but they also have relevant applications in areas such as nonlinear elastic mechanics \cite{Zhikov2}, electrorheological fluids \cite{Ruzicka} or image restoration \cite{LiLiPi}. The reader can find detailed discussions of the properties of variable Lebesgue and Sobolev spaces in the recent books \cite{Cruz,Die} and the review papers \cite{di}, \cite{sam}, as well as the references therein.

There are also other types of spaces with variable exponents that have been studied. For example, Hardy spaces and Campanato spaces with variable exponents are discussed in \cite{Nakai}. Besov spaces with variable exponents are investigated in \cite{AlHa2010}. The variable exponent Besov space $B^{\alpha(\cdot)}_{p(\cdot),q(\cdot)}$ is defined by using the mixed Lebesgue-sequence space $\ell^{q(\cdot)} (L^{p(\cdot)})$. Some properties of the space $\ell^{q(\cdot)} (L^{p(\cdot)})$ are derived in \cite{AlHa2010, GG,GRS, KeVy}.

In this paper, we explore some geometric properties of the space $\ci$, such as strictly and uniformly convexity as well as the approximation identity, by using the property of reflexivity that was proved in \cite{GRS}.

Let us  revisit the notion of variable exponent Lebesgue spaces. We denote by $\mathcal{P}_0$ the set of measurable functions $p(\cdot): \mathbb{R}^n \rightarrow [c,\infty]$, where $c>0$, and by $\mathcal{P}$ the subset of $\mathcal{P}_0$ such that the range of its elements is contained in $[1,\infty]$.

We define $p_{+}= \mathop{\rm ess \; sup}\limits p$ and $p_{-}= \mathop{\rm ess \; inf}\limits p$ as the essential supremum and infimum of the variable exponent $p(\cdot)\in \mathcal{P}_0$. Let $\Omega_{\infty}=\{x\in \mathbb{R}^n: p(x)=\infty\}$ and $\Omega_0=\mathbb{R}^n\setminus \Omega_{\infty}$ be the subsets of $\mathbb{R}^n$ where the exponent is infinite and finite, respectively. The variable exponent Lebesgue space $L^{p(\cdot)}(\mathbb{R}^n)$ consists of all measurable functions $f$ such that there exists a positive constant $\lambda$ for which the modular

\[\varrho_{L^{p(\cdot)}}(\lambda^{-1}f)
:=\int_{\Omega_0}\left(\frac{|f(x)|}{\lambda}\right)^{p(x)}
dx
+\left\|\lambda^{-1}f\right\|_{L^{\infty}(\Omega_{\infty})}
\]
is finite. Given a variable exponent  $p(\cdot)\in \mathcal{P}_0$   and a function $f \in L^{p(\cdot)}(\mathbb{R}^n)$, we introduce the Luxemburg norm associated with the variable exponent Lebesgue space as
\[
\|f\|_{L^{p(\cdot)}}=\inf\{\lambda>0:  \varrho_{L^{p(\cdot)}}(\lambda^{-1}f)\leq 1\}.
\]

Next, we introduce the notion of the variable mixed spaces
$\ell^{q(\cdot)} (L^{p(\cdot)}(\mathbb{R}^n))$
following the definition in \cite{AlHa2010}. Let $p(\cdot), q(\cdot) \in \mathcal{P}_0$
and $(f_{\nu})_{\nu}$ be a sequence of measurable functions such that $f_\nu\in L^{p(\cdot)}(\mathbb{R}^n)$ for every $ \nu\in\mathbb{N} $. The modular of the variable mixed space is given by
\begin{align*}\label{P1}
\varrho_{{\ell}^{q(\cdot)}(L^{p(\cdot)})}\Big((f_{\nu})_{\nu}\Big)&:= \sum_{\nu=1}^{\infty}\inf\Big\{\lambda_{\nu}>0 : \varrho_{p(\cdot)}\left(\frac{f_{\nu}}{\lambda_{\nu}^{\frac{1}{q(\cdot)}}}\right)\le 1\Big\},
\end{align*}
where we adopt the convention $\lambda^{\frac{1}{\infty}} = 1$. When $ q_+<\infty $, this modular can also be expressed as
\begin{equation*}\label{P1a}
\varrho_{{\ell}^{q(\cdot)}(L^{p(\cdot)})}\Big((f_{\nu})_{\nu}\Big)=\underset{\nu}{\sum}\|f_{\nu}^{q(\cdot)}\|_{\frac{p(\cdot)}{q(\cdot)}}.
\end{equation*}
The variable mixed space  $\ell^{q(\cdot)} (L^{p(\cdot)}(\mathbb{R}^n))$ is then defined as
\begin{equation*}\label{P2}
{\ell}^{q(\cdot)}(L^{p(\cdot)}):=\Big\{f=(f_\upsilon)_\upsilon |~\exists\lambda>0~:~~
\varrho_{{\ell}^{q(\cdot)}(L^{p(\cdot)})}\left((f_{\upsilon})_{\upsilon}/\lambda\right)<\infty \Big\}.
\end{equation*}
If $p(\cdot), q(\cdot) \in \mathcal{P}_0$, then the space $\ell^{q(\cdot)} (L^{p(\cdot)}(\mathbb{R}^n))$ is
a quasi-normed space, i.e.,
\begin{align*}
\|f\|_{{\ell}^{q(\cdot)}(L^{p(\cdot)})}= \inf\left\{\mu>0 : \varrho_{{\ell}^{q(\cdot)}(L^{p(\cdot)})}\left(f/\mu\right)\le 1\right\},
\end{align*}
which is a quasi-norm on  $\ell^{q(\cdot)} (L^{p(\cdot)}(\mathbb{R}^n))$.
As shown in \cite[Theorem 1]{KeVy} and \cite[Theorem 3.6]{AlHa2010},
$\|\cdot\|_{{\ell}^{q(\cdot)}(L^{p(\cdot)})}$ becomes a norm, if  $p(\cdot), q(\cdot) \in \mathcal{P}$ satisfy one of the following conditions:
\begin{enumerate}
\item
$1 \le q(\cdot) \le p(\cdot) \le \infty $,
\item
$p(\cdot)\ge 1$
and $q(\cdot) \ge 1$ is constant,
\item
$\frac{1}{p(\cdot)} + \frac{1}{q(\cdot)}\leq 1 $.
\end{enumerate}
Often we need to assume some additional regularity on $ p(\cdot) $. There are two important continuity condition as follows \cite{Cruz},
\begin{itemize}
\item
We say that $ p(\cdot) $ is locally log-H\"{o}lder, and we write $ p(\cdot)\in LH_0(\mathbb{R}^n) $, if
\begin{align*}
\exists C_0\quad:\quad |x-y|<\frac{1}{2}\longrightarrow~|p(x)-p(y)|\leq\frac{C_0}{-\log (|x-y|)}
\end{align*}

\item
We say that $ p(\cdot) $ is log-H\"{o}lder continuous at infinity, and we write $ p(\cdot)\in LH_\infty(\mathbb{R}^n) $, if
\begin{align*}
\exists C_0,~\exists p_\infty\quad:\quad \forall x\in\mathbb{R}^n\longrightarrow~|p(x)-p_\infty|\leq\frac{C_\infty}{\log (e+|x|)}
\end{align*}
where $ p_\infty=\lim_{x\rightarrow\infty}p(x) $.
\end{itemize}
We will denote $ p(\cdot)\in LH(\mathbb{R}^n) $, if $ p(\cdot) $ is log-H\"{o}lder continuous locally and at infinity. Throughout this paper, inequalities for exponents are understood in the almost everywhere sense. Our results are true as long as the space $ \ci $ is a Banach space.

\begin{definition}\label{defi:220130}
 For a vector-valued sequence space $S(X)$
 over a reflexive Banach space $X,$ define its K\"othe dual with
respect to the dual pair $(X, X^*)$ (see
\cite{GKP}) as follows:
\[
S(X)':=\left\{\bar{\varphi}=(\varphi_\nu)_{\nu=1}^\infty \in {X^{*}}^{\mathbb{N}}: \mbox{for each}~~ \bar{f}
=(f_\nu)_{\nu=1}^\infty \in S(X), \sum_{\nu=1}^\infty |\varphi_\nu(f_\nu)|<\infty \right\}.
\]
The norm of $\bar{\varphi}=(\varphi_\nu)_{\nu=1}^\infty \in S(X)'$ is given by
\[
\|\bar{\varphi}\|_{S(X)'}:=
\sup\left\{
\sum_{\nu=1}^\infty |\varphi_\nu(f_\nu)|\,:\,\bar{f}
=(f_\nu)_{\nu=1}^\infty \in S(X),~
\|\bar{f}\|_{S(X)} \le 1
\right\}.
\]
\end{definition}

We recall that there is a canonical isomorphism between the dual and the K\"{o}the dual of space $\ci$. Indeed, by \cite[Lemma 2.4]{GRS} and \cite[Definition 2.2]{GRS}, the mapping
\begin{align}\label{Kothedual}
\psi:\ci'&\rightarrow(\ci)^*
\\
f'&\mapsto\psi_{f'}\qquad \text{s.t.}\quad\|f'\|_{\ci'}=\|\psi_{f'}\|_{(\ci)^*}\notag
\end{align}
is a canonical isomorphism, where $ \ci' $ denotes the K\"{o}the dual of space $\ci$, and for every $ g\in\ci $ we have
\begin{align*}
\psi_{f'}(g)=\sum_{\nu=1}^{\infty}\int_{\mathbb{R}}f_\nu'g_\nu dx\qquad:\qquad f_\nu'\in L^{p'(\cdot)}, g_\nu\in L^{p(\cdot)}
\end{align*}
where $ p'(\cdot) $ is conjugate exponent of $ p(\cdot) $.

To conclude, we give the definition of the maximal function and revisit its established boundedness within the variable exponent Lebesgue space \( L^{p(\cdot)} \).

\begin{definition}
Recall that the Hardy-Littlewood maximal operator $ \mathcal{M} $ is defined on $ L^1_{loc} $ by
\begin{align*}
\mathcal{M}f(x)=\sup_{r>0}\frac{1}{|B(x,r)|}\int_{B(x,r)}|f(y)|~dy
\end{align*}
where $ B(x, r) $ denotes the ball with center $ x\in\mathbb{R}^n $ and radius $ r > 0 $.
\end{definition}

In the last section, we need to invoke the following theorem
on the boundedness of the maximal operator in $L^{p(\cdot)}$;
see \cite[Theorem 3.16]{Cruz}.
\begin{theorem}\label{T1}
Given a set $\Omega$ and $p(\cdot) \in \mathcal{P}(\Omega),$ if $\frac{1}{p(\cdot)} \in \mbox{LH}(\Omega),$ then
$$
\|\mathcal{M} f\|_{L^{p(\cdot)}(\Omega)} \le C \, \|f\|_{L^{p(\cdot)}(\Omega)}
$$
for all $f \in L^{p(\cdot)}(\Omega)$.
\end{theorem}

\section{Strict Convexity}

In this section, we demonstrate the strict convexity of the space $\ci$ using its reflexivity property. The space $\ci$ is a mixed sequence space that has a complex structure and poses some challenges for working with it. We begin by recalling some basic definitions and the facts that are crucial for proving our result.

\begin{definition}\cite[Page 101]{LP}
A Banach space $ X $ is called strictly convex if the unit ball $ B(X) $ of $ X $ satisfies the following property: for any two distinct vectors $ x,y \in B(X) $,  $ |x + y| < 2 $.
\end{definition}

\begin{definition}\cite{LP}\label{def:Smooth}
A Banach space $ X $ is said to be Gateaux smooth (or smooth) if every point $ x \in S(X) $ on the unit sphere $ S(X) $ of $ X $ is smooth, i.e., for any $ y\in X $, the limit
\begin{align*}
\lim_{t\rightarrow 0}\frac{\|x+ty\|_X-\|x\|_X}{t}
\end{align*}
exists.
\end{definition}

\begin{definition}
Let $ X $ be a Banach space. We define the set-valued mapping $ J $ on $ S(X) $ as follows,
\begin{align*}
J(x)=\{x^*\in S(X^*):~\langle x^*,x\rangle=1 \}.
\end{align*}
\end{definition}
By the Hahn-Banach theorem, for any $ x\in S(X) $, the set $ J(x) $ is nonempty.

\begin{remark}\label{J:reflexivity}
Let $ 1<q_-,p_-,q_+,p_+<\infty $, then $ \ci $ is reflexive \cite[Theorem 1.1]{GRS}. The following mapping is a canonical isomorphism,
\begin{align*}
\Gamma:\ci &\longrightarrow(\ci)^{**}
\\
& f\mapsto\Gamma_{f}\qquad s.t.\qquad \|f\|_{\ci}=\|\Gamma_f\|.
\end{align*}
Then, for $ f^*\in (\ci)^* $, we have
\begin{align*}
J(f^*)&=\{f\in S((\ci)^{**}): \langle f,f^*\rangle =1\}
\\
&=\{~f\in S(\ci)~\quad: \langle f,f^*\rangle =1\}\subset S(\ci).
\end{align*}
\end{remark}

\begin{definition}\cite{LP}
Let $ X $ be a Banach space. The mapping $ J $ is said to be norm to $ weak^* $ continuous at $ x $, if $ (x_n)_n\subseteq S(X) $ such that $ lim_{n\rightarrow\infty} x_n = x $, and $ x_n^*\in J(x_n) $ for all $ n\in\mathbb{N} $, then
$ (x_n^*)_n $ converges $ weak^* $ to an element in $ J(x) $.
\end{definition}

\begin{theorem}\cite[Theorem 2.2.4]{LP}\label{theo:SC,Smooth}
Let $ X $ be a Banach space.
\begin{itemize}
\item[(1)] If $ X^* $ is strictly convex, then $ X $ is smooth.
\item[(2)] If $ X^* $ is smooth, then $ X $ is strictly convex.
\end{itemize}
\end{theorem}

The main result of this section is the following theorem, which shows that the space $\ci$ is strictly convex, by using the Helly's theorem.

\begin{theorem}\label{theo:smooth}
Let $ 1<q_-,p_-,q_+,p_+<\infty $, then $ (\ci)^* $ is smooth.
\end{theorem}
\begin{proof}
According to \cite[Theorem 2.2.2]{LP}, it suffices to show that at every point $ f^*\in(\ci)^* $, the mapping $ J $ is a norm to $ weak^* $ continuous at $ f^* $.

Let $ (f_n^*)_n\subseteq S((\ci)^*):\lim_{n\rightarrow\infty}f_{n}^*=f^* $ and $ f_n\in J(f_n^*) $ for all $ n\in\mathbb{N} $. On the other side, since $ \ci $ is a reflexive Banach space \cite{GRS} and it is also separable \cite{GG}, then $ (\ci)^* $ is separable too.
 
Therefore, $ (\ci)^* $ is separable and
by Remark \ref{J:reflexivity}. $ (f_n)_n\subseteq S(\ci) $ is a bounded sequence in $ (\ci)^{**}\approx\ci $, then by Helly's theorem \cite[Theorem 27.11]{Helly'sTheorem}, it has a $ weak^* $ convergent subsequence like $ (f_{n_i})_{n_i} $. Then,
\begin{align*}
\exists f\in \ci\quad s.t.\quad f_{n_i}\overset{w^*}{\longrightarrow }f,
\end{align*}
by considering the reflexivity of the space,
\begin{align*}
 f_{n_i}\overset{w}{\longrightarrow }f,
\end{align*}
then
\begin{align*}
\|f\|_{\ci}\leq\limsup_{n_i\to\infty}\|f_{n_i}\|_{\ci}=1.
\end{align*}
We have, $ \|f_{n_i}^*-f^*\|_{(\ci)^*}\rightarrow 0 $ then 
\begin{align*}
\|f^*\|_{(\ci)^*}=1,
\end{align*}
 and by \eqref{Kothedual} for $ f\in S(\ci) $,
\begin{align*}
f_{n_i}\overset{w}{\longrightarrow }f\Leftrightarrow~\langle f,f^*\rangle &=\lim_{n_i\rightarrow\infty}\langle f_{n_i},f^*\rangle \qquad:\qquad f^*\in(\ci)^*
\\
&=\lim_{n_i\rightarrow\infty}\langle f_{n_i},f'\rangle ~\qquad:\qquad f'\in\ci'
\\
&=\lim_{n_i\rightarrow\infty}\langle f_{n_i},f_{n_i}'\rangle \qquad:\qquad\|f_{n_i}'-f'\|_{\ci'}\rightarrow 0
\\
&=\lim_{n_i\rightarrow\infty}\langle f_{n_i},f_{n_i}^*\rangle \qquad:\qquad \langle f_{n_i},f_{n_i}^*\rangle =1
\\
&~=\lim_{n_i\rightarrow\infty} 1=1.
\end{align*}
then
\begin{align*}
\|f\|_{\ci}=\|\Gamma_f\|=\sup_{\|f^*\|_{(\ci)^*}\leq 1}\langle f,f^*\rangle =1.
\end{align*}
Therefore,
\begin{align*}
 f\in S(\ci)\quad:\quad \langle f,f^*\rangle =1,
\end{align*}
then $ f\in J(f^*) $, which is the desired result.
\end{proof}
\begin{corollary}
Let $ 1<q_-,p_-,q_+,p_+<\infty $, then $ \ci $ is strictly convex.
\end{corollary}
\begin{proof}
Since $ \ci $ is a Banach space, then by Theorem \ref{theo:SC,Smooth} and Theorem \ref{theo:smooth} we have the desired result.
\end{proof}

\section{Uniform Convexity}

Here, we examine if the variable mixed Lebesgue-sequence space $ \ell^{q(\cdot)}(L^{p(\cdot)}) $ is uniformly convex. We find that uniform convexity is not always a feature of this space, especially when $ q=\infty $ or $ q=1 $.

Let's review the concepts associated with uniform convexity.

\begin{definition}\label{defUC}
A Banach space $X$ is said to be uniformly convex if for all $\varepsilon\in(0,2]$ there exists $\delta>0$ such that
\[
\|\frac{x+y}{2}\|\le 1-\delta
\]
whenever $ x,y \in S(X) $ satisfies $\|x-y\|\ge \varepsilon$.
\end{definition}
\begin{definition}
A normed vector space X is said to be uniformly smooth if every $ \varepsilon>0 $ there exists $ \delta>0 $ such that if $ x,y\in X $ with $ \|x\|_{X}=1 $ and $ \|y\|_{X}\leq\delta $ then
\[\|x+y\|_{X}+\|x-y\|_{X}\leq 2+\varepsilon\|y\|_{X}.\]
\end{definition}

Recall that a Banach space $ X $ is uniformly smooth if for any $x$ in unit sphere $ S(X) $ and $ y $ in $ X $, the limit
\begin{align}\label{T2}
\lim_{t\to 0}\Phi_{x,y}(t):=\lim\limits_{t \to 0} \frac{\|x-ty\|_{X}-\|x\|_{X}}{t}
\end{align}
exists and is uniform in $x, y \in S(X)$  \cite{LP}.

In the following proposition, we will see that the question of whether the space is uniformly convex can have either a positive or negative answer, depending on specific cases.
\begin{prop}\label{propNotUC}
Let $ p(\cdot)\in\mathcal{P} $. The following statements are true.
\begin{itemize}
\item[1.] If $q \in (1,\infty)$ is a constant and $1<p_{-} \le p_{+} < +\infty$, then the space ${\ell}^{q}(L^{p(\cdot)})$ is uniformly convex.
\item[2.] The space $\ell^{\infty}(L^{p(\cdot)})$ is not uniformly convex.
\item[3.] The space $ \ell^1(L^{p(\cdot)}) $ is not uniformly convex.
\end{itemize}
\end{prop}
\begin{proof}
We prove each statement separately.
\begin{itemize}
\item[1.]
By \cite[Proposition 3.3]{AlHa2010}, we have
\[
\|(f_{\mu})_{\mu}\|_{\ci}=\|\|f_{\mu}\|_{\pe}\|_{\ell^q}.
\]
Since $L^{p(\cdot)}$ is uniformly convex, by \cite[Theorem 3]{Day} it follows that ${\ell}^{q}(L^{p(\cdot)}(\Omega))$ is also uniformly convex.

\item[2.]
To show that $\ell^{\infty}(L^{p(\cdot)})$ is not uniformly convex, it suffices to consider the functions $f=(\chi_{(0,1)}(\cdot), \chi_{(0,1)}(\cdot), \cdots )$ and $g=(0, \chi_{(0,1)}(\cdot),0,\chi_{(0,1)}(\cdot), \cdots )$.  Then we have
\begin{align*}
\|f\|_{\ell^{\infty}(L^{p(\cdot)})} &= \|g\|_{\ell^{\infty}(L^{p(\cdot)})}=1 \\
\|f-g\|_{\ell^{\infty}(L^{p(\cdot)})}=1 &~~~~\mbox{and}~~~~ \|\frac{f+g}{2}\|_{\ell^{\infty}(L^{p(\cdot)})}=1,
\end{align*}
which is the desired result.

\item[3.] Similarly, to show that $ \ell^1(L^{p(\cdot)}) $ is not uniformly convex, we can consider the functions $f=(\chi_{(0,1)}(\cdot), 0,0, \cdots )$ and $g=( 0,\chi_{(0,1)}(\cdot),0,0, \cdots )$. Then we have
\begin{align*}
  \|f\|_{\ell^{1}(L^{p(\cdot)})} &= \|g\|_{\ell^{1}(L^{p(\cdot)})}=1,   \\
 \quad \|f-g\|_{\ell^{1}(L^{p(\cdot)})}=2 &~~~~\mbox{and}~~~~ \|\frac{f+g}{2}\|_{\ell^{1}(L^{p(\cdot)})}=1
\end{align*}
which also violates the definition of uniform convexity.
\end{itemize}
This completes the proof.
\end{proof}

The following theorem aids in examining cases where the exponent $q$ is variable.

\begin{theorem}\label{T1}
\cite[Theorem 2.2.5]{LP} For every Banach space, $ X $ is uniformly convex if and only if its dual space $ X^* $ is uniformly smooth.
\end{theorem}

Initially, we demonstrate that the quantity $ \Phi_{f,g}(t) $ is non-decreasing as a function of $t$.
\begin{lemma}\label{rem:nondecreasing}
Let $ f,g\in\ci $, then $ \Phi_{f,g}(t) $ is  non-decreasing on $ t\in\mathbb{R}^+ $.
\end{lemma}
\begin{proof}
According to existing isomorphic \eqref{Kothedual}, we have
\begin{align*}
\exists  f',g'\in \ci'  \quad:\quad \psi_{f'}=f,~ \psi_{g'}=g.
\end{align*}
Define $ \varphi_{f',g'}(t):=\frac{\langle f'-tg',h\rangle-1}{t} $ for $ \|h\|_{\ci}\leq 1 $. Then one can see $\varphi_{f',g'}(t)$ is a non-decreasing mapping on $ \mathbb{R}^+ $ i.e.  for every $ 0<t_1\leq t_2 $,
\[
\varphi_{f',g'}(t_1)\leq \varphi_{f',g'}(t_2).
\]
We note that in smooth space definition we assume 
\[
~\|f'\|_{\ci'}=\|\psi_{f'}\|_{(\ci)^*}=\|f\|_{(\ci)^*}~= 1. 
\]
Then for any $\|h\|_{\ci}\leq 1$, we have 
\[
\langle f',h\rangle~\leq~1.
\]
In sequentially, for $ 0< t_1\leq t_2 $, one can write 
  
\begin{align*}
&\qquad\varphi_{f',g'}(t_1)\leq \varphi_{f',g'}(t_2)
\\
\Leftrightarrow&
\qquad
\frac{\langle f'-t_1g',h\rangle -1}{t_1}\leq~ \frac{\langle f'-t_2g',h\rangle -1}{t_2}
\\
\Leftrightarrow&
\qquad
\langle f'-t_1g',t_2h\rangle -t_2\leq~ \langle f'-t_2g',t_1h\rangle -t_1
\\
\Leftrightarrow&
\qquad
\langle f'-t_1g',t_2h\rangle -\langle f'-t_2g',t_1h\rangle ~\leq~t_2-t_1
\\
\Leftrightarrow&
\qquad
\langle f',(t_2-t_1)h\rangle -t_1t_2\langle g',h\rangle +t_1t_2\langle g',h\rangle ~\leq~t_2-t_1
\\
\Leftrightarrow&
\qquad
\langle f',(t_2-t_1)h\rangle ~\leq~t_2-t_1
\\
\Leftrightarrow&
\qquad
\langle f',h\rangle ~\leq~1
\end{align*}
Now, by taking supremum over $ \|h\|_{\ci}\leq 1 $, one can write that the mapping $ \Phi_{f,g}(t) $ is also  nondecreasing with respect to $ t\in\mathbb{R}^+ $. 
\end{proof}

We are now equipped to address the question of the uniform convexity of $\ci$. 
\begin{theorem}\label{theoUC}
Let $ 1<q_-.p_-,q_+,p_+<\infty $, then
$ \ci $ is uniformly convex.
\end{theorem}
\begin{proof}
Thanks to Theorem \ref{theo:smooth}. and Definition \ref{def:Smooth}. the limit in \eqref{T2} exists. Then one can write, for every $ f,g\in(\ci)^* $, there exists a $ c(f,g) $ such that
\begin{align*}
\lim_{t\rightarrow 0}\Phi_{f,g}(t)=c(f,g),
\end{align*} 
in other words, for every $ f,g\in(\ci)^* $ we have
\begin{align}\label{Pointwiselimit}
\forall\varepsilon>0,~\exists\delta_1(\varepsilon,f,g):~t\in B(0,\delta_1(\varepsilon,f,g))
\longrightarrow~ c(f,g)-\frac{\varepsilon}{2}<\Phi_{f,g}(t)<\frac{\varepsilon}{2}+c(f,g).
\end{align}
Now, we are going to show that the limit in \eqref{T2} is uniform in $ f,g\in S(X^*) $ where $ X:=\ci $. We need to show that
\begin{align*}
\lim_{t\rightarrow 0}\sup_{f,g\in S(X^*)}|\Phi_{f,g}(t)-c(f,g)|
=0,
\end{align*}
in other words,
\begin{align*}
\forall\varepsilon>0,\quad\exists\delta(\varepsilon)\quad:~t\in B(0,\delta(\varepsilon))~\longrightarrow~\sup_{f,g\in S(X^*)}|\Phi_{f,g}(t)-c(f,g)|<\varepsilon.
\end{align*}
By redutio and absurdum argument,
\begin{align}\label{reductioandabsurdum}
& \exists\varepsilon_0>0,\quad\forall\delta(\varepsilon_0)~:\quad\exists t_0\in B(0,\delta(\varepsilon_0))~
\\
\&\quad&\sup_{f,g\in S(X^*)}|\Phi_{f,g}(t_0)-c(f,g)|\geq\varepsilon_0>\frac{\varepsilon_0}{2},\nonumber
\end{align}
then $ \exists f_0,g_0\in S(X^*) $ such that
\begin{align*}
|\Phi_{f_0,g_0}(t_0)-c(f_0,g_0)|>\frac{\varepsilon_0}{2}.
\end{align*}
Let us deal with following two cases:
\begin{itemize}
\item[Case 1:] If $ t_0\in[0,\delta(\varepsilon_0)] $: By Remark \ref{rem:nondecreasing} we have
\begin{align*}
\frac{\varepsilon_0}{2}&< |\Phi_{f_0,g_0}(t_0)-c(f_0,g_0)|
\\
&=|\Phi_{f_0,g_0}(t_0)-\inf_{0<t}\Phi_{f_0,g_0}(t)|
\\
&=\Phi_{f_0,g_0}(t_0)-\inf_{0<t}\Phi_{f_0,g_0}(t)
\\
&=\Phi_{f_0,g_0}(t_0)-c(f_0,g_0), 
\end{align*}
then by \eqref{Pointwiselimit} and \eqref{reductioandabsurdum} and considering $ \varepsilon:=\varepsilon_0,~ t:=t_0 $ and $ \delta(\varepsilon_0):=\delta_1(\varepsilon_0,f_0,g_0) $, we have
\begin{align*}
\exists\varepsilon_0,\quad\exists\delta(\varepsilon_0),\quad\exists t_0\in [0,\delta(\varepsilon_0)]\quad\&\quad
c(f_0,g_0)+\frac{\varepsilon_0}{2}<\Phi_{f_0,g_0}(t_0)<\frac{\varepsilon_0}{2}+c(f_0,g_0),
\end{align*}
which is a contradiction.
\item[Case 2:] If $ t_0\in[-\delta(\varepsilon_0),0) $: then $ -t_0\in (0,\delta(\varepsilon_0)] $ and 
\begin{align*}
-t\in[0,\delta(\varepsilon_0)]~:~\Phi_{f,g}(-t)=-\Phi_{f,-g}(t)~\Longrightarrow~\lim_{t\rightarrow 0}\Phi_{f,g}(-t)=-c(f,-g),
\end{align*}
and by Remark \ref{rem:nondecreasing} we have
\begin{align*}
\frac{\varepsilon_0}{2}&
<|\Phi_{f_0,g_0}(-t_0)-\inf_{0<-t}\Phi_{f_0,g_0}(-t)|
\\
&=\Phi_{f_0,g_0}(-t_0)-\inf_{0<-t}\Phi_{f_0,g_0}(-t)
\\
&= \Phi_{f_0,g_0}(-t_0)-\lim_{t\rightarrow 0}\Phi_{f_0,g_0}(-t)
\\
&=\Phi_{f_0,g_0}(-t_0)+c(f_0,-g_0),
\end{align*}
thus as in the Case 1, one can write
\begin{align*}
\exists\varepsilon_0,~\exists\delta(\varepsilon_0),~\exists -t_0\in [0,\delta(\varepsilon_0)]~\&~
-c(f_0,-g_0)+\frac{\varepsilon_0}{2}<\Phi_{f_0,g_0}(-t_0)<\frac{\varepsilon_0}{2}-c(f_0,-g_0),
\end{align*}
which is a contradiction.
\end{itemize}
Therefore, as $ t\rightarrow 0 $, the limit in \eqref{T2} exists and is uniform in $ f,g\in S((\ci)^*) $. Thus by Theorem \ref{T1} and relation \eqref{T2}, the desired result is obtained.
\end{proof}

\section{Convergence in measure}
Now, we aim to establish that, given certain conditions on the exponents $ \p, \q\in\mathcal{P} $, the mixed variable Lebesgue spaces exhibit the property where convergence in norm also entails convergence in measure.

During the proof of the subsequent theorem, similar to the approach in \cite{GRS}, we use $P_N$ to represent the projection onto $U_N$. Here, $U_N$ is defined as the set containing all sequences $(f_\nu)_{\nu=1}^\infty$ for which $f_\nu=0$ whenever $\nu > N$.

\begin{theorem}\label{NormImpliesMeauser}
Suppose that $ \p,\q \in\mathcal{P} $ and $ q_+,p_+<\infty, $ then the space $ \ci $ has the property that convergence in norm implies convergence in measure.
\end{theorem}

\begin{proof}
Let $ \lambda\in(0,1) $ and $ A_N(n):=\{x:|P_N(f_n-f)(x)|>\lambda \} $, then 
\begin{align*}
A_N(n)&=\{x:|P_N(f_n-f)(x)|>\lambda \}
\\
&=\{x:\Big(\sum_{\nu=1} ^N|(f_{\nu_n}-f_\nu)(x)|^2\Big)^{\frac{1}{2}}>\lambda \}
\\
&\subseteq\{x:\sum_{\nu=1}^N|(f_{\nu_n}-f_\nu)(x)|>\lambda \}
\\
&\subseteq\underset{\nu=1}{\overset{N}{\bigcup}}~\{x:|(f_{\nu_n}-f_\nu)(x)|>\frac{\lambda}{N} \}.
\end{align*}
We have $ \|f_n-f\|_{\ci}\rightarrow 0 $, then we without loss of generality, may consider $ \|f_n-f\|_{\ci}\leq 1 $. For every $ N\in\mathbb{N} $ we have
\begin{align*}
|A_N(n)|&=|\{x:|P_N(f_n-f)(x)|>\lambda \}|
\\
&\leq \sum_{\nu=1}^N|\{x: |(f_{\nu_n}-f_\nu)(x)|>\frac{\lambda}{N}\}|
\\
&\leq(\frac{N}{\lambda})^{p_+} \sum_{\nu=1}^N\int|(f_{\nu_n}-f_\nu)(x)|^{p(x)} dx
\\
&=(\frac{N}{\lambda})^{p_+}\sum_{\nu=1}^N\varrho_{\frac{p(\cdot)}{q(\cdot)}}(|f_{\nu_n}-f_\nu|^{q(\cdot)})
\\
&\leq(\frac{N}{\lambda})^{p_+}\sum_{\nu=1}^N\||f_{\nu_n}-f_\nu|^{q(\cdot)}\|_{\frac{p(\cdot)}{q(\cdot)}}
\\
&\leq(\frac{N}{\lambda})^{p_+}\varrho_{\ci}(f_n-f)
\\
&\leq(\frac{N}{\lambda})^{p_+}\|f_n-f\|_{\ci}.
\end{align*}
Therefore, for every $ N\in\mathbb{N} $, we have $ \underset{n\rightarrow\infty}{\lim}|A_N(n)|=0 $. In other words,
\begin{align}\label{limofn}
\forall N,\quad\forall\varepsilon>0,\quad\exists\delta'(N,\varepsilon):\quad n\geq\delta'(N,\varepsilon)\longrightarrow |A_N(n)|<\frac{\varepsilon}{2}.
\end{align}
Let $ A(n):=\{x:|(f_n-f)(x)|>\lambda \} $, then
\begin{align*}
A(n)&=\{x:|(f_n-f)(x)|>\lambda \}
\\
&=\{x:\lim_{N\to\infty}|P_N(f_n-f)(x)|>\lambda \}
\\
&=\underset{N=1}{\overset{\infty}{\bigcup}}\{x:|P_N(f_n-f)(x)|>\lambda \}
\\
&=\underset{N=1}{\overset{\infty}{\bigcup}} A_N(n).
\end{align*}
Since, for every $ n $, the sequence $ (A_N(n))_N $ is a nondecreasing sequence
\begin{align}\label{sup of sequs}
|A(n)|&=|\underset{N=1}{\overset{\infty}{\bigcup}} A_N(n)|\nonumber
\\
&=\lim_{N\to\infty}|A_N(n)|
\\
&=\sup_{N\in\mathbb{N}}|A_N(n)|.\nonumber
\end{align}
Now we are going to show that
\begin{align*}
\lim_{n\rightarrow\infty}|A(n)|=0.
\end{align*}
By reductio and absurdum, we have
\begin{align*}
\exists\varepsilon_0,\quad\forall \delta,\quad\exists n_0\geq\delta\quad\&\quad |A(n_0)|\geq\varepsilon_0>\frac{\varepsilon_0}{2},
\end{align*}
and by \eqref{sup of sequs}, one can write
\begin{align}\label{reductio and absurdum}
\exists\varepsilon_0,\quad\forall \delta,\quad\exists n_0\geq\delta\quad\&\quad\exists N_0:\quad |A_{N_0}(n_0)|>\frac{\varepsilon_0}{2}.
\end{align}
Let in \eqref{limofn}, the values $ N:=N_0,~\varepsilon:=\varepsilon_0 $, then in the \eqref{sup of sequs} set $ \delta:=\delta'(N_0,\varepsilon_0) $. Thus we have
\begin{align*}
\exists n_0\geq\delta: \quad\frac{\varepsilon_0}{2}<|A_{N_0}(n_0)|<\frac{\varepsilon_0}{2},
\end{align*}
which is a contradiction, and this completes the proof.
\end{proof}

\section{Approximate Identity}

Following the methodology outlined in the \cite{Cruz}, we apply the technique of approximate identities, also referred to as mollification, within the space $ \ell^1(L^\frac{\p}{\q}) $. For a given function $ \phi $ and each $ t > 0 $, we define $\phi_t(x)=t^{-n}\phi(x\slash t) $. This normalization ensures that if $ \phi $ belongs to $ L^1 $, then the $ L^1 $ norm of $ \phi_t $ remains consistent with that of $ \phi $. The radial majorant of $ \phi $ is then determined by the function
\begin{align*}
\Phi(x)=\sup_{|y|>|x|}\phi(y).
\end{align*}

\begin{definition}
For a function $ \phi $ in $ L^1 $ with the property that $ \int_{\mathbb{R}^n}\phi(x)dx=1 $, the collection $ \{\phi_t\}=\{\phi_t:~t>0 \} $ is known as an approximation identity. When the radial majorant of $ \phi $ belongs to $ L^1 $ as well, this collection $ {\phi_t} $ is referred to as a potential type approximation identity.
\end{definition}

If $ f $ belongs to $ L^p $ for $ p $ in the range $ [1,\infty] $ and $ \{\phi_t\} $ constitutes a potential type approximation identity, then it follows that
\begin{align*}
\phi_t * f(x) \to f(x),
\end{align*}
almost everywhere for each point, as $ t $ approaches zero.

In the following Theorem \ref{theo2.10}, we aim to demonstrate that the space $ \ell^1(L^{\p\slash\q}) $ upholds the approximation identity.  It is important to note that achieving such a result is not possible in a general setting.

\begin{theorem}\label{theo1}
Let $ \p,\q \in\mathcal{P} $ such that  $ q_+,p_+<\infty, $ and $ \{T_t\}_{t} $ be a family of linear operators and $ \mu $ as a measure on $ \mathbb{R}^n $. We will define $ T^*f(x):=\underset{t}{\sup}|T_tf(x)| $ for  $ x\in\mathbb{R}^n. $ Suppose that there exist
$(d,c,\alpha)\in(\mathbb{R}^{\geq 0})^3$ such that
\begin{equation}\label{firstCondition}
\mu\{x\in\mathbb{R}^n:~|T^*f(x)|>\lambda \}\leq\Big(\frac{c\|f\|_{\ci}}{\lambda^\alpha}\Big)^d,
\end{equation}
for all $\lambda>0$ and
\begin{equation}\label{secondCondition}
\lim_{t\rightarrow t_0}T_tP_Nf(x)=P_Nf(x),
\end{equation}
almost everywhere for all $N \in {\mathbb N}$. Then the set $ \{ f\in\ci:~\underset{t\rightarrow t_0}{\lim}T_tf(x)=f(x)~a.e.\} $  is closed in $ \ci $.
\end{theorem}
\begin{proof}
Let $ \|f\|_{\ci}\leq 1,~ \|f-P_Nf\|_{\ci}\longrightarrow 0 $ and  for every $ N\in \mathbb{N}$ we have $ \underset{t\rightarrow t_0}{\lim}T_tP_Nf(x)=P_Nf(x) $ almost everywhere.  Then by the linearity of $ T_t $ and the definition of $ T^* $, for every $ \lambda>0, $ we have
\begin{align*}
&~\quad\mu\{x:~\limsup_{t\rightarrow t_0}|T_tf(x)-f(x)|>\lambda \}
\\
&\leq\mu\{x:~|T^*(f-P_Nf)(x)|>\frac{\lambda}{2} \}+\mu\{x:~|(f-P_Nf)(x)|>\frac{\lambda}{2} \}
\\
&\leq \Big(\frac{c~\|f-P_Nf\|_{\ci}}{(\lambda\slash 2)^\alpha}\Big)^d~+\mu\{x:~|(f-P_Nf)(x)|>\frac{\lambda}{2} \}.
\end{align*}
By Theorem \ref{NormImpliesMeauser}, we have 
\begin{align*}
\mu\{x:~|(f-P_Nf)(x)|>\frac{\lambda}{2} \}\longrightarrow 0,
\end{align*} 
thus
\begin{align}\label{astast}
\mu\{x:~\limsup_{t\rightarrow t_0}|T_tf(x)-f(x)|>\lambda \}=0
\end{align}
for all $\lambda>0$. In addition,
\begin{align*}
\{x:~\limsup_{t\rightarrow t_0}&|T_tf(x)-f(x)|>0 \}\subseteq\underset{k=1}{\bigcup^\infty}\{x:~\limsup_{t\rightarrow t_0}|T_tf(x)-f(x)|>\frac{1}{k} \}.
\end{align*}
Therefore, by \eqref{astast} we have
\begin{align*}
\mu\{x:~\limsup_{t\rightarrow t_0}|T_tf(x)-f(x)|>0 \}
&\leq\mu\Big(\underset{k=1}{\bigcup^\infty}\{x:~\limsup_{t\rightarrow t_0}|T_tf(x)-f(x)|>\frac{1}{k} \}\Big)
\\
&\leq\sum_{k=1}^{\infty}\mu\{x:~\limsup_{t\rightarrow t_0}|T_tf(x)-f(x)|>\frac{1}{k}\}
=0,
\end{align*}
which is the desired result.
\end{proof}

We observe that for $ \p,\q\in\mathcal{P} $ and $ q_+,p_+<\infty $, the condition $ f\in L^{p(\cdot)} $ is equivalent to
$ f^{q(\cdot)}\in L^{\frac{p(\cdot)}{q(\cdot)}} $. Hence, the following definition is well-posed.

\begin{definition}\label{def1}
Let $ \p,\q\in\mathcal{P} $ and $ q_+,p_+<\infty $, then define
\begin{align*}
V:=\Big\{g=(f_\nu^\q)_\nu~\Big|~(f_\nu)_\nu\in(L^{p(\cdot)})^{\mathbb{N}}~~ \&~~\|g\|_V=\|(f_\nu^{\q})_\nu\|_{\ell^1(L^{\p\slash\q})}<\infty\Big\}.
\end{align*}
\end{definition}
We note that the set $ V\subset\underset{\nu\in\mathbb{N}}{\prod} L^{\frac{\p}{\q}} $ is well-defined.

\begin{lemma}\label{lem1}
Let $ \p,\q \in\mathcal{P} $ and $ q_+,p_+<\infty, $ and $ \phi $ be a positive radial decreasing and integrable function. Then for $ g\in V $ we have
\begin{align*}
\underset{t}{\sup}|(\phi_t*g)(x)|\leq\|\phi\|_1\underset{\nu=1}{\sum^\infty} \mathcal{M}f_\nu^{q(\cdot)}(x).
\end{align*}
\end{lemma}
\begin{proof}
Recall that for every  $ \nu $ we have $ \Big[\underset{t}{\sup}|(\phi_t*f_\nu^\q)(x)|\Big]^2\leq\Big[\|\phi\|_1\mathcal{M}f_\nu^\q(x)\Big]^2. $ Then
\begin{align*}
\sup_t|(\phi_t*g)(x)|&\leq \Big(\sum_{\nu=1}^\infty\Big[\sup_t|(\phi_t*f_\nu^\q)(x)|\Big]^2 \Big)^\frac{1}{2}
\\
&\leq\|\phi\|_1 \Big(\sum_{\nu=1}^\infty\Big[\mathcal{M}f_\nu(x)\Big]^2 \Big)^\frac{1}{2}
\\
&\leq\|\phi\|_1 \sum_{\nu=1}^\infty\mathcal{M}f_\nu(x),
\end{align*}
which is the desired result.
\end{proof}

\begin{lemma}\label{lem2}
Let $ \p,\q \in\mathcal{P} $ where $ q_+,p_+<\infty, $ and $ 1<(\frac{p}{q})_-,~\frac{\p}{\q}\in LH(\mathbb{R}^n)$ and $ \mu $ as a measure on $ \mathbb{R}^n $. Then we can find $(d,c,\alpha)\in(\mathbb{R}^+)^3$ such that 
\begin{align*}
\mu\{x\in\mathbb{R}^n:~\sum_{\nu=1}^\infty\mathcal{M}f_\nu^{q(\cdot)}(x)>\lambda \}\leq\Big(\frac{c~\|(f_\nu^{\q})_\nu\|_{\ell^1(L^{p(\cdot)/q(\cdot)})}}{\lambda^\alpha}\Big)^{d}.
\end{align*}
for all $\lambda>0$.
\end{lemma}
\begin{proof}
Let $ \|(f_\nu^{\q})_\nu\|_{\ell^1(L^{p(\cdot)/q(\cdot)})}\leq 1 $. Since $ 1<(\frac{p}{q})_-,~\frac{\p}{\q}\in LH(\mathbb{R}^n), $ then by \cite[Proposition 2.3]{Cruz} and \cite[Theorem 3.16]{Cruz}, there exists a constant $ c_1 $ depending on the dimension $ n $, the log-H\"older
constants of $ \frac{\p}{\q},~ (\frac{p}{q})_- $ and $ (\frac{p}{q})_\infty $ (if this value is finite), such that
\begin{align*}
\|\sum_{\nu}\mathcal{M}f_\nu^{\q}\|_{\frac{p(\cdot)}{q(\cdot)}}
&\leq\sum_{\nu}\|\mathcal{M}f_\nu^{\q}\|_{\frac{p(\cdot)}{q(\cdot)}}
\\
&\leq c_1\sum_{\nu}\|f_\nu^{\q}\|_{\frac{p(\cdot)}{q(\cdot)}}
\\
&\leq c_1~\|(f_\nu^{\q})_\nu\|_{\ell^1(L^{p(\cdot)/q(\cdot)})}
\\
&\leq c_1.
\end{align*}
Thus,
\begin{align}\label{eq1}
\varrho_{\frac{p(\cdot)}{q(\cdot)}}(\frac{\sum_{\nu}\mathcal{M}f_\nu^{\q}}{c_1\sum_{\nu}\|f_\nu^{\q}\|_{\frac{p(\cdot)}{q(\cdot)}}})\leq 1.
\end{align}
Let $ \lambda\in(0,1) $. Without loss of generality, we may consider  $ c_1\sum_{\nu}\|f_\nu^{\q}\|_{\frac{p(\cdot)}{q(\cdot)}}\geq 1 $. Therefore by \eqref{eq1} we have
\begin{align*}
\varrho_{\frac{p(\cdot)}{q(\cdot)}}\Big(\sum_{\nu}\mathcal{M}f_\nu^{\q}\Big)
&\leq\Big(c_1\sum_{\nu}\|f_\nu^{\q}\|_{\frac{p(\cdot)}{q(\cdot)}}\Big)^{(\frac{p}{q})_+}
\\
&=\Big(c_1\|(f_\nu^{\q})_\nu\|_{\ell^1(L^{p(\cdot)/q(\cdot)})}\Big)^{(\frac{p}{q})_+}
\\
&\leq c_1^{(\frac{p}{q})_+}\|(f_\nu^{\q})_\nu\|_{\ell^1(L^{p(\cdot)/q(\cdot)})},
\end{align*}
thus,
\begin{align*}
\mu\{x:~\sum_{\nu}\mathcal{M}f_\nu^{q(\cdot)}(x)>\lambda \}&\leq\int_{\mathbb{R}^n}\Big(\frac{\sum_{\nu}\mathcal{M}f_\nu^{q(\cdot)}(x)}{\lambda}\Big)^{\frac{p(x)}{q(x)}}d\mu
\\
&\leq~\frac{1}{\lambda^{(\frac{p}{q})_+}}\int_{\mathbb{R}^n}\Big(\sum_{\nu}\mathcal{M}f_\nu^{q(\cdot)}(x)\Big)^{\frac{p(x)}{q(x)}}d\mu
\\
&=~\frac{1}{\lambda^{(\frac{p}{q})_+}}~\varrho_{\frac{\p}{\q}}\Big(\sum_{\nu}\mathcal{M}f_\nu^\q\Big)
\\ 
&\leq \frac{c_1^{(\frac{p}{q})_+}\|(f_\nu^{\q})_\nu\|_{\ell^1(L^{p(\cdot)/q(\cdot)})}}{\lambda^{(\frac{p}{q})_+}}.
\end{align*}
Now by considering $ c=c_1^{(\frac{p}{q})_+},~\alpha=(\frac{p}{q})_+ $ and $ d=1 $, we have the desired result.
\end{proof}
\begin{corollary}\label{coro3}
Let $ g \in V $. Due to Lemmas \ref{lem1} and \ref{lem2}, the operator 
\begin{align*}
g\mapsto\sup_{t}|\phi_t*g|,
\end{align*}
satisfies the condition \eqref{firstCondition}.
\end{corollary}
\begin{lemma}\label{coro4}
Let $ \p\in\mathcal{P} $ and $ p_+<\infty, $ then the condition \eqref{secondCondition} is satisfied for $ V $.
\end{lemma}
\begin{proof}
Recall that since $ p_+<\infty $ then for $ f_\nu\in L^{\p} $ we have $ f_\nu^{\q}\in L^{\frac{\p}{\q}} $, therefore, by \cite[Theorem 5.8]{Cruz}, for $ g\in V $ we have
\begin{align*}
\lim_{t\rightarrow 0}~\phi_t*P_Ng(x)=P_Ng(x),
\end{align*}
almost everywhere for all $N \in {\mathbb N}$.
\end{proof}
\begin{theorem}\label{theo2.10}
Let $ \p,\q \in\mathcal{P} $ and $ q_+,p_+<\infty, $ and $ 1<(\frac{p}{q})_-,~\frac{\p}{\q}\in LH(\mathbb{R}^n)$. Then the approximation identity holds in V.
\end{theorem}
\begin{proof}
As in \cite[Lemma 2.1]{GRS} we have $ \bigcup\limits_{N=1}^\infty (V \cap U_N) $ is dense in $ V $. If we combine this fact with Corollary \ref{coro3}, Lemma \ref{coro4} and  Theorem \ref{theo1}, then we will have the desired result.
\end{proof}
\begin{corollary}
Let $ \p\in\mathcal{P},~ p_+<\infty $ and $ 1<p_-,~\p\in LH(\mathbb{R}^n) $. Then the approximation identity holds in $ \ell^1(L^\p) $.
\end{corollary}
\begin{proof}
According to the Theorem \ref{theo2.10}, it suffices to consider $ \q=1 $ almost everywhere.
\end{proof}

\bibliographystyle{amsplain}

\end{document}